### v1

# Foreground Removal in Ground-Based CMB Observations Using a Transformer Model

Abstract
We present a novel method for Cosmic Microwave Background (CMB) foreground removal based on deep learning techniques. This method employs a Transformer model, referred to as TCMB, which is specifically designed to effectively process HEALPix-format spherical sky maps. TCMB represents an innovative application in CMB data analysis, as it is an image-based technique that has rarely been utilized in this field. Using simulated data with noise levels representative of current ground-based CMB polarization observations, the TCMB method demonstrates robust performance in removing foreground contamination. The mean absolute variance for the reconstruction of the noisy CMB Q/U map is significantly less than the CMB polarization signal. To mitigate biases caused by instrumental noise, a cross-correlation approach using two half-mission maps was employed, successfully recovering CMB EE and BB power spectra that align closely with the true values, and these results validate the effectiveness of the TCMB method. Compared to the previously employed convolutional neural network (CNN)-based approach, the TCMB method offers two significant advantages: (1) It demonstrates superior effectiveness in reconstructing CMB polarization maps, outperforming CNN-based methods. (2) It can directly process HEALPix spherical sky maps without requiring rectangular region division, a step necessary for CNN-based approaches that often introduces uncertainties such as boundary effects. This study highlights the potential of Transformer-based models as a powerful tool for CMB data analysis, offering a substantial improvement over traditional CNN-based techniques.
[…]
1 Introduction
The Cosmic Microwave Background (CMB) provides crucial observational information for precision cosmology (e.g., Bennett et al., 2003; Story et al., 2013; Das et al., 2014; Planck Collaboration et al., 2016). The next focus of research lies in the precise measurement of CMB polarization. To this end, the scientific community has deployed more advanced polarization detection experiments, such as ground-based telescope arrays like the Simons Observatory (SO)(Ade et al., 2019), BICEP Array (Hui et al., 2018), CMB-S4 (Abazajian et al., 2016), Ali CMB Polarization Telescope (AliCPT, Li et al., 2018), as well as space-based missions like LiteBIRD (Matsumura et al., 2014), CORE (Delabrouille et al., 2018), and PICO (Hanany et al., 2019). These CMB experiments are designed to achieve multifrequency coverage and high sensitivity. This progress undoubtedly poses greater challenges for data processing, making the exploration of innovative methods to break through traditional data processing techniques essential.
One of the major contaminants in CMB observations is Galactic diffuse foregrounds, which require specialized techniques to remove from the observed data. For CMB polarization, the primary components of these foregrounds are synchrotron and thermal dust emissions, which contribute polarized signals at levels several orders of magnitude higher than the CMB signal across almost all observed frequency bands (BICEP2/Keck Collaboration et al., 2015; Krachmalnicoff et al., 2016). Various approaches are employed to analyze multifrequency CMB observations, aiming to separate the CMB signal from foreground contamination. Component separation methods can be broadly categorized into two types: blind methods and non-blind methods. Non-blind methods typically rely on prior knowledge of the emission characteristics of foregrounds. For example, the Commander method (Eriksen et al., 2008) utilizes Gibbs sampling to estimate the joint foreground-CMB posterior distribution. In contrast, the ForeGroundBuster method (Stompor et al., 2009) employs an analytically derived likelihood function for spectral parameters. However, a key limitation of these methods is their dependence on prior knowledge of foregrounds. When faced with unknown or complex foregrounds, such approaches may introduce biases in the detection of CMB B-modes (Armitage-Caplan et al., 2012; Remazeilles et al., 2016; Hensley & Bull, 2018). In contrast, blind source component separation algorithms do not rely on prior knowledge of foregrounds. Instead, they leverage the independent statistical properties of different components to separate signals. Examples of blind methods include spectral matching independent component analysis (Delabrouille et al., 2003), independent component analysis (Baccigalupi et al., 2004), and variants of the internal linear combination (ILC) method (ILC, Eriksen et al., 2004; Tegmark et al., 2003) (NILC, Delabrouille et al., 2009).
Focusing on foreground contamination removal, deep learning methodologies have been extensively studied in recent years, with Convolutional Neural Networks (CNNs) demonstrating significant potential for reconstructing CMB signals (Wang et al., 2022; Casas et al., 2022a; Yan et al., 2023a; Casas et al., 2023; Yan et al., 2024; Sudevan & Chen, 2024; Pal et al., 2024; Adak, 2025). Moreover, these methods have been widely extended to other CMB-related tasks, such as CMB lensing reconstruction (Caldeira et al., 2019; Yan et al., 2023c; Heinrich et al., 2024; Flöss et al., 2024), CMB delensing (Yan et al., 2023b; Heinrich et al., 2024), and point source processing (Bonavera et al., 2021; Casas et al., 2022b; Lambaga et al., 2025). However, CNNs are inherently limited to processing flat images within Euclidean space, necessitating the transformation of HEALPix (Górski et al., 2005) spherical sky maps into flat projections. This transformation inevitably introduces artificial boundaries, which can compromise data integrity. In fact, deep learning methodologies capable of effectively processing HEALPix spherical sky maps remain an active area of research. Krachmalnicoff & Tomasi (2019) extended the one-dimensional CNN approach to HEALPix maps by reorganizing pixels according to HEALPix pixelization properties and performing one-dimensional convolutions. While innovative, this method introduces additional computational complexity. In contrast, DeepSphere (Perraudin et al., 2019) proposed a deep learning model using Graph Convolutional Networks (GCNs) tailored for HEALPix sky maps, treating each pixel as a node in a graph structure. However, the large number of pixels in standard observational sky maps imposes significant computational burdens on graph networks. Studies by Petroff et al. (2020) and Adams et al. (2023) indicate that GCNs tend to underperform compared to CNNs in foreground removal for CMB data.
The Transformer model (Vaswani et al., 2017), as explored by Liu et al. (2021) and Carlsson et al. (2023), shows promise in processing HEALPix sky maps. However, these studies primarily focus on spherical images encountered in daily life and have not been extensively applied to CMB maps. In this paper, we introduce a novel method based on the Transformer model, referred to as TCMB (Transformer for Cosmic Microwave Background). We utilize this method to remove foreground contamination from CMB polarization data in ground-based observations. Our results demonstrate that TCMB can effectively process HEALPix sky maps and handle regions of arbitrary size, offering a robust and flexible solution for CMB data analysis.
The paper is organized as follows. In Section 2, we describe the mock data. Section 3 provides a comprehensive introduction to the TCMB method. The results of foreground removal are presented in Section 4. Finally, we summarize and conclude in Section 5.
2 Mock Data Sets
Deep learning methods rely heavily on large datasets to optimize models effectively. In this section, we describe the process for generating the dataset utilized in this study.
We adopt a ground-based CMB polarization experiment as the baseline for generating our dataset. Specifically, we use the AliCPT-like experiment (Li et al., 2018; Ghosh et al., 2022; Dou et al., 2024; Zhang et al., 2024b, a) as our instrument configuration baseline, as it is one of the telescopes currently in operation and is capable of hosting large detector arrays to provide highly precise measurements of CMB polarization in the short near future. The AliCPT-1 experiment aims to measure CMB polarization signals at two frequencies, 95 GHz and 150 GHz. The observational sky patch for AliCPT-1 is located in the Northern Hemisphere, centered at R.A. = 180° and decl. = 30°, covering approximately 17% of the sky.
Figure 1 shows the observational sky patch of AliCPT-1 using the “20 K” mask. This mask is generated by excluding pixels with a noise standard deviation above “20 K” at =1024 for the 150 GHz channel, resulting in an accepted sky fraction of about 10%. The “UNP” mask further removes the sky above a declination of 65° to mitigate foreground contamination, reducing the accepted sky fraction to approximately 6.7%. In this work, we adopt the “UNP” mask. In addition to the data from the 95 GHz and 150 GHz bands of AliCPT-1, we combine data from the Planck experiment to construct multi-frequency datasets, by simulating data in the Planck High Frequency Instrument (HFI)’s four bands: 100 GHz, 143 GHz, 217 GHz, and 353 GHz.
The CMB simulation are Gaussian realizations generated from the power spectra. Specifically, we initially employ the CAMB111https://github.com/cmbant/CAMB (Lewis et al., 2000) software package to compute the CMB power spectra and lensing power spectrum within the framework of cold dark matter cosmology (CDM). The tensor-to-scalar ratio is fixed to . Other cosmological parameters relevant to the CDM framework include , , , , , and . The best-fit value and corresponding standard deviations for these parameters are derived from the Planck 2018 parameters (Planck Collaboration et al., 2020). The lensing effects are added by the LensPyx (Reinecke et al., 2023) package.
In the foreground simulation, we take into account various components of the Galactic diffuse foreground polarization, including thermal dust, synchrotron radiation, and anomalous microwave emission (AME). The foreground maps are generated from the PySM3 222https://github.com/galsci/pysm/tree/main package (Thorne et al., 2017; Zonca et al., 2021). The simulation of synchrotron radiation is performed by extrapolating template maps based on a parametric model derived from the PySM s1 model. The thermal dust and AME derived from the PySM d1 model and a2 model.
Experiment | AliCPT | Planck HFI |
Frequency (GHz) | 95 150 | 100 143 217 353 |
Beam size (arcmin) | 19 150 | 9.7 7.3 5.0 4.9 |
The instrumental noises for AliCPT-1 are Gaussian realizations generated from the noise standard deviation maps of AliCPT-1 bands. The noise of AliCPT-1 exhibits anisotropy, with the distribution of noise levels presented in Figure 1. This noise level is derived from four years of observational data and the the total integrated observation time is projected to be “48 module*years”. The harmonic mean of the noise level are 3.2 K-arcmin for 95 GHz and 4.9 K-arcmin for 150 GHz. Here, we utilize the half-mission noise level, which is enhanced by a factor of in comparison to the full-mission noise level. Planck half-mission noise maps are based on the FFP10 noise simulations and downloaded from the Planck Legacy Archive (PLA) 333https://www.cosmos.esa.int/web/planck/pla. For the half-mission noise simulation, PLA provides only 300 noise realizations, resulting in a total of 600 noise maps when considering two half-mission. To augment the number of noise maps, we employ a linear combination method. The specific steps are as follows: (1) Generate 10 random numbers () to calculate the corresponding weights: ; (2) Randomly select 10 noise maps from the total of 600 noise maps; (3) Multiply these 10 maps by their respective weights and sum them to obtain a new noise map.
The simulated maps for each channel include CMB and foregrounds, all smoothed with the instrumental Gaussian beams as listed in Table 1. In total, we generated fifty thousand observed sky maps with NSIDE=512. For each realization, the cosmological parameters were randomly sampled within a range of five standard deviations. Additionally, the spectral index maps and amplitude maps of the foreground models were multiplied by a random number drawn from a Gaussian distribution with a mean of 1 and a standard deviation of 0.1 (0.05 for the spectral index). This approach is designed to account for uncertainties in cosmological parameters as well as the complexity of the foreground model. Further details can be found in reference (Yan et al., 2024).
3 Method
Transformer is a relatively novel neural network architecture that primarily employs the self-attention mechanism (Bahdanau et al., 2014) to extract intrinsic features (Vaswani et al., 2017), demonstrating significant potential for widespread application in artificial intelligence (AI). Transformer was initially implemented in the domain of natural language processing (NLP), where it demonstrated substantial enhancements in performance (Vaswani et al., 2017; Devlin et al., 2018). Brown et al. (2020) pre-trained a large transformer-based model known as GPT-3 (short for Generative Pre-trained Transformer 3), which comprises 175 billion parameters. It demonstrated impressive performance across various natural language processing tasks. Inspired by the significant success of transformer architectures in the field of NLP, researchers have applied this architecture to computer vision (CV) tasks. The ViT model (Dosovitskiy et al., 2020) is the first implementation of the Transformer architecture in the domain of computer vision (CV). In the field of CV, CNNs have long been regarded as fundamental components; however, the transformer architecture is now emerging as a potential alternative to CNNs. To date, the Transformer has demonstrated its efficiency and has shown strong competitive performance compared to CNNs. In this work, we utilize the Transformer to remove the polarized foregrounds of CMB.
3.1 Patches
The standard Transformer receives as input a 1D sequence of tokens. To handle 2D images, the input image in ViT model is initially segmented, flattened, and subsequently transformed into lower-dimensional linear representations referred to as patch embeddings. Specifically, image is reshaped into a sequence of flattened 2D patches , where (, ) is the resolution of the original image, C is the number of channels, (, ) is the resolution of each image patch, and is the resulting number of patches. These flattened patches are mapped to a dimensional space through a trainable linear projection layer, and can be represented in matrix form as . Thus, we convert image into a sequence that can be used as input for the Transformer, with a sequence length of . Inspired by the ViT model, we transform the HEALPix map into sequence data suitable for processing by Transformer, based on the pixelization characteristics of the HEALPix representation.
HEALPix (Calabretta, 2004; Górski et al., 2005) is a general structure for the pixelization of data on the sphere. It offers an authentic curvilinear partitioning of the sphere into quadrilaterals of exactly equal area, distinguished by their varied geometrical shapes. Specifically, HEALPix partitions the sphere into equal-area grids, where NSIDE represents the resolution of the grid. The boundaries for NSIDE=1 define the base-resolution pixels and higher-order pixelisations are defined by their regular subdivision. NSIDE describes the number of subdivisions along the edges of a base-resolution pixel required to achieve the desired high-resolution partition. HEALPix support two different numbering schemes for the pixels, RING scheme and NESTED scheme. The former is suited for spherical harmonic analysis. The latter refers to a pixel nesting arrangement that conveniently elucidates the structural configuration of the pixels and the relationships among adjacent pixels. In the sky map of the NESTED scheme, the inherent nesting pattern of pixels facilitates a natural division of the sky map into multiple equal-area patches. Inspired by the ViT model, we aggregate these patches into a sequence data suitable for processing by Transformer architecture.
We consider a full-sky map with resolution parameter NSIDE. First, we define the parameter to determine the size of the patch, where needs to be a power of two and NSIDE. Subsequently, the entire full-sky map in NESTED scheme can be naturally segmented into multiple patches, with the total number of patches being and the number of pixels in each patch being . Figure 2 illustrates an example in which the full-sky map with NSIDE=8 is partitioned into multiple patches, with a parameter of . In the context of analyzing a partial sky, one can employ a mask to identify and retain patches that correspond to the observational coverage, while excluding those that are outside this coverage. This methodology facilitates a significant reduction in the number of patches and, as a result, optimizes the use of computational resources. Finally, we arrange the patches according to their latitude, thereby obtaining sequence data suitable for processing by the Transformer.
Given a full-sky map , it can be partitioned into multiple patch sequence data based on its nested characteristics. Here, , Npix, and represent the number of patches, and pixels in the full-sky map and pixels in a patch, respectively. If we consider the multi-band sky maps, denoted as , then the corresponding patch sequence can be represented as , where represent the number of frequency bands. Subsequently, we apply a linear projection layer (or one-dimensional convolution layer) to generate the input data suitable for processing by the Transformer, where is a constant. The parameters involved in this operation () are acquired through learning. The aim of this operation is to enable the transformation of map patches into vector representations, referred to as patch embeddings (tokens), thereby enabling the Transformer model to process map as sequences of tokens using a transformer-based methodology.
3.2 Self-attention mechanism
The core component of the Transformer architecture is the self-attention (SA) mechanism, which can determine the importance of a single token (patch embedding) relative to other tokens in the sequence by computing the relevance between that token and the other tokens (Vaswani et al., 2017). This mechanism effectively models long-range dependencies, which has resulted in the widespread recognition of the SA mechanism as a key contributor to the success of Transformer models.
Suppose the input data is a set of tokens , where and is the number of tokens and token dimension. The objective of SA is to capture the interactions among all tokens by encoding each token in relation to the global contextual information. That is done by defining three learnable weight matrices () to transform input into Queries (), Keys (), and Values ():
where and are the hidden dimension. The corresponding attention matrix can be expressed as follows:
where softmax is normalized exponential function. Then the SA is performed as follows
where is the output of SA.
The limited capacity of a single SA module often leads to its focus on only a few positions, potentially overlooking other important positions. In practice, the implementation of multi-head SA (MSA) is predominant. MSA employs the parallel stacking of SA blocks (heads), thereby enhancing the efficacy of the self-attention layer. The idea of MSA is to stack multiple SA blocks in parallel, and the concatenated outputs are fused by a projection matrix . Each head has its own learnable weight matrices denoted by , where and denotes total number of heads in MSA block. The output of -th head SA is given by formula 3. The output of MSA can write,
3.3 Network architecture
The network architecture employed in this work is illustrated in Figure 3. As outlined in Section 3.1, the masked multiband observation maps are initially partitioned into multiple patches. Subsequently, these patches are mapped to patch embedding sequence through a linear layer, which are then ultimately input into the Transformer module.
We use Swin Transformer block (STB) (Liu et al., 2021) as the fundamental unit of Transformer module. Two successive STBs are presented in the panel (c) of Figure 3. Each STB is comprised of a LayerNorm (LN) layer, a MSA module, a residual connection, and a two-layer multi-layer perceptron (MLP) employing GELU non-linearity. Different from the conventional MSA module, the first block uses window-based MSA (W-MSA), and the second block shifts those windows and applies window-based MSA again (SW-MSA). W-MSA initially partitions tokens into multiple windows, with each window containing several non-overlapping tokens and performing SA calculations independently. Since the number of windows is significantly smaller than the total number of tokens, W-MSA effectively reduces computational complexity compared to MSA. However, there is no interaction between different windows. To facilitate information exchange between distinct windows, the second block employs the SW-MSA module. SW-MSA adjusts the window allocation through a sliding mechanism, ensuring that the newly defined windows contain tokens that differ from those in the first block. After the reallocation of windows, self-attention is performed independently on each window once again. Consequently, the two STBs not only significantly alleviate computational burdens but also enable information exchange among all tokens, thereby enhancing computational efficiency. Further detailed information can be found in reference (Liu et al., 2021).
The design of the Transformer architecture is based on reference (Liang et al., 2021), consisting of four layers of residual swin transformer block (RSTB), with each RSTB layer composed of three sets of two successive STBs. The output of the Transformer is subsequently transformed through a linear mapping to obtain the target sky map.
3.4 Dataset and Training
In accordance with the description in Section 2, we simulated fifty thousand sets of multi-band sky maps with NSIDE=512 to construct the training set. Additionally, we independently simulated two thousand sets of sky maps for the test set. The Planck noise data in the test set was obtained by downloading it from the PLA platform. The inputs to the network model consist of beam-convolved observational Q or U maps derived from all six frequency bands. The desired output of the network is a beam-convolved CMB Q or U map with beam size of 95 GHz, plus noise at 95 GHz. In reference (Yan et al., 2024), we highlighted the challenges faced by deep learning methods in effectively separating the CMB polarization from noise on the map level. Therefore, in this study, our training objective continues to take into account data that includes both the CMB and noise, while concentrating the network model’s efforts on the removal of the foreground. We select the noise in the 95 GHz frequency band as the desired output noise of the network, primarily due to the lower noise level associated with this band compared to others. A detailed discussion on the selection of frequency bands can be found in reference (Yan et al., 2024).
First, we preprocess the observational sky maps, utilizing a method derived from the ILC. The ILC method requires the observed data to be deconvolved to a common beam and resolution (Delabrouille et al., 2009; Remazeilles et al., 2011). Here, we employ this method for preprocessing the observed data. We consider multi-frequency observational sky maps , wherein each frequency band is characterized by distinct instrumental beam profiles. The index and denote frequency band and pixel, respectively. The observed maps convolves/deconvolutes to the same resolution in harmonic space:
where, is harmonic coefficients of . and denote the beam window function associated with each frequency band and the common beam window function, respectively. In order to maintain a correspondence with the desired output of the network, the observational maps are also reconvolved with a common beam of the 95 GHz. However, it is important to note that in our tests, we found that this data preprocessing method did not have a significant impact on the results compared to the case without preprocessing. If there are substantial errors in the instrumental beam, this preprocessing may be deemed unnecessary.
After the pre-processing of the multi-band observed sky maps, it is necessary to partition the maps into multiple patches. The sky map has NSIDE=512, and we employ a parameter of for the partitioning. Consequently, the full-sky map can be divided into patches, with each patch containing 64 pixels. In our tests, slight variations in , such as increasing it to or decreasing it to , have a negligible impact on the results. We utilize a mask to retain the patches covered by observations while excluding those that are not observed. The effective number of patches is 3494. To facilitate the computations in the Transformer model, we employ zero padding to expand the number of patches to 3600. This procedure is solely intended to facilitate computation, and the padded patches will not be considered in subsequent data analyses.
During the training process of the neural network model, the optimization of model parameters is achieved by minimizing the loss function. The loss function serves to quantify the discrepancy between the network’s predictions and the target. In this study, we have chosen to employ the mean absolute deviation (L1) loss function as our optimization criterion. To optimize network model’s performance, we employ the Adam optimizer, as proposed by Kingma & Ba (2014), initializing the learning rate at and gradually decreasing it over the course of the iterations until it reaches a final value of . The training process consists of approximately 200 epochs, with a batch size of 24, and is conducted on eight NVIDIA RTX 4090 GPUs. On average, the training process for a single network model takes approximately 70 hours to complete.
4 Results
In this section, we employ the proposed method, TCMB, to conduct foreground removal on CMB polarization data and present the outcomes of this analysis.
4.1 Recovered CMB map
We first present the results of foreground removal at the map level. As illustrated in Figure 4, the recovered CMB polarization map, including noise components, shows a high degree of consistency with the true map. The information remaining in the residual map is minimal, indicating that the diffuse foreground has been effectively and cleanly removed. To quantitate the effectiveness of foreground removal, we employed the calculation of the mean absolute deviation (MAD):
where is the number of pixels, and represent the predicted and real sky maps. The learning objective of the network model is to minimize the MAD value as much as possible, aiming for it to approach zero. In the test set, the values of the MAD are K for recovery of CMB Q map and K for recovery of CMB U map. These values are approximately 1/200 of the CMB polarization signal, indicating that the residual foreground is quite limited. This further validates the high efficiency of our method in effectively removing foreground contamination.
4.2 Recovered CMB power spectra
After obtaining the foreground-removed CMB map, we utilized the NaMaster package (Alonso et al., 2019) to compute the power spectrum. The function, as implemented in NaMaster, was utilized to apodize the UNP mask prior to the computation of the power spectra, employing apodization scales of . Figure 5 illustrates the power spectrum of the CMB map reconstructed by the network model, which includes instrumental noise. The error bars are derived from statistical results obtained from the test set and represent the standard deviation between the power spectrum calculated from the target CMB map and that computed from the CMB map reconstructed by the network model.
For the E-mode power spectrum, we can see that the CMB EE spectrum reconstructed by the network model is in high agreement with the target power spectrum. However, due to the presence of instrumental noise in the recovered CMB map, the power spectrum exhibits a significant deviation from the true CMB EE power spectrum at smaller scales. The BB power spectrum is predominantly dominated by instrumental noise. We can see that the BB power spectrum reconstructed by the network model is also in strong agreement with the target power spectrum. These results indicate that the diffuse foreground can be effectively removed, thereby preserving the CMB and its associated noise information in full.
To obtain a cleaner CMB power spectra, it is necessary to apply debiasing techniques of noise to the recovered CMB maps by network. Here, we employ two half-mission maps for cross-correlation analysis to facilitate the debiasing of noise. We can segment the observed time-ordered data (TOD) from the entire survey mission into two halves, corresponding to two half-mission datasets. These two data subsets contain the same CMB signal and foregrounds, but their noise is uncorrelated. By calculating the cross-correlation power spectrum between the two half-mission maps, we can effectively reduce the noise present in the power spectrum. We utilize the network model to perform individual foreground removal on each half-mission map, resulting in two foreground-cleaned CMB sky maps. Subsequently, we calculate their cross-correlation power spectrum.
The power spectra after noise debiasing is presented in Figure 6. The error bars represent the statistical uncertainty derived from the test set, with this uncertainty primarily originating from instrument noise. Notably, at smaller scales, the uncertainty introduced by instrument noise becomes significantly pronounced. We can see that the reconstructed CMB EE and BB power spectra are consistent with the true values, indicating that TCMB foreground cleaning pipelines is highly effective.
4.3 Comparison with CNN method
References (Yan et al., 2023a, 2024) have demonstrated a method for CMB foreground removal based on convolutional neural networks (CNNs), known as CMBFSCNN. Here, we present a comparison between the results of the CMBFSCNN method and those of the TCMB method.
The CNN method is limited to processing flat images, necessitating the conversion of HEALPix spherical sky maps into flat sky maps. As illustrated in Figure 7, the observed sky area is artificially divided into two independent regions within the flat image, introducing non-natural boundaries. In our implementation, we convert the spherical sky maps from the dataset into flat sky maps and use them to train the CMBFSCNN network model. The output maps from the CMBFSCNN network consist of beam-convolved CMB maps combined with noise at 95 GHz.
The results of foreground removal are illustrated in Figure 8. For clarity, we only present the residual maps of the recovered CMB Q and U, which help highlight the differences between the two methods. Compared to Figure 4, the residual maps show significantly more pronounced artifacts, indicating that the CMBFSCNN method is less effective than the TCMB method in recovering the noisy CMB Q and U maps. To facilitate a quantitative comparison, we calculated the MAD values between the reconstructed CMB maps and the target sky maps. For the test set, the MAD values are K for recovery of CMB Q map and K for recovery of CMB U map. These MAD values are approximately three times larger than those obtained using the TCMB method, and they demonstrate that, at the map level, the TCMB method is more efficient than the CMBFSCNN method for foreground removal.
Furthermore, it is noteworthy that the residual map in Figure 8 exhibits two prominent boundary marks, which arise from the artificially introduced boundaries during the transformation of HEALPix spherical sky maps into flat maps. These additional boundaries disrupt the spatial continuity of the signal, leading to contamination at the discontinuous boundaries during the foreground removal process and resulting in edge effects. This highlights a significant drawback of using the CNN method for processing HEALPix spherical sky maps. In contrast, the Transformer-based TCMB method can directly process HEALPix spherical sky maps, effectively mitigating edge effects. As a result, the TCMB method demonstrates a clear advantage over the CNN method when handling HEALPix spherical sky maps.
The power spectra of the noisy CMB maps reconstructed using the CMBFSCNN network are presented in Figure 8. For comparison, the results obtained using the TCMB method are also included in the figure. We observe that the noisy CMB EE and BB power spectra reconstructed by the CMBFSCNN network are in substantial agreement with the target spectra; however, minor biases are present, particularly at smaller scales where these biases are more pronounced. It is important to note that these discrepancies remain relatively small compared to the noisy CMB power spectra. In contrast, the CMB power spectrum obtained using the TCMB method exhibits negligible bias. Furthermore, the uncertainties in the power spectrum recovered by the TCMB method are significantly smaller than those obtained using the CMBFSCNN method across all angular scales. Therefore, from the perspective of the power spectrum, the TCMB method demonstrates a significant advantage over the CMBFSCNN method in the reconstruction of noisy CMB maps.
We employ cross-correlation analysis of two half-mission maps to eliminate noise biases in CMB power spectra. The results are presented in Figure 10. We can see that the CMB EE and BB power spectra obtained using the CMBFSCNN method are consistent with the true values. In this analysis, no significant advantage of the TCMB method over the CMBFSCNN method in reconstructing the CMB power spectra was identified. This is primarily attributed to the fact that the main contribution to the errors shown in the Figure 10 originates from instrument noise. While the TCMB method demonstrates superior performance in reconstructing noisy CMB Q/U maps compared to the CMBFSCNN method, this advantage becomes significantly obscured after the removal of noise biases due to the uncertainties introduced by noise. Consequently, the noise level plays a dominant role in the uncertainty associated with the recovery of the CMB polarization signals. Additionally, during the process of debiasing the power spectrum, the network model does not directly participate, as it primarily processes the sky map data. A more equitable comparison should be based on recovered maps and noisy power spectra as shown Figures 9. Therefore, the conclusion that the TCMB method outperforms the CMBFSCNN method in removing CMB foregrounds remains valid. The Transformer model employed by the TCMB method has the primary disadvantage of requiring significantly more computational resources during the training process compared to CNN method.
4.4 Dependency of foreground models
In this section, we investigate the dependency of the network model on the foreground model. In the data from the training set, synchrotron radiation is simulated based on the s1 model from the PySM package, while thermal dust radiation is simulated using the d1 model. After the network has completed its training, we perform a re-simulation of the test dataset. In the test set data, synchrotron radiation is simulated with the s2 model, and thermal dust radiation is based on the d4 model. Compared to the s1 model, the spectral index of the s2 model varies with the galactic latitude. Additionally, the d1 model represents a single-component thermal dust radiation, whereas the d4 model incorporates two components of thermal dust. Consequently, the foreground model in the test set differs from that in the training set. We input the test set data into the fully trained network model to assess the model’s dependency on the foreground model.
The residual maps between the recovered CMB maps by TCMB and the target maps are presented in Figure 11. Compared to the results in Figure 4, we can see that the residual maps contain more information. This indicates that the alteration of the foreground model in the test dataset has led to a reduction in the efficiency of the network in foreground removal, resulting in a small amount of foreground remaining that could not be effectively removed. To quantitatively assess the results, we also calculated the MAD value. In the test set for TCMB, the values of the MAD are K for recovery of CMB Q map and K for recovery of CMB U map. Due to the presence of a small amount of residual foreground, these MAD values are slightly greater compared to the results presented in Figure 4. We also employed the CMBFSCNN method to conduct a same assessment of the dependency on the foreground model. After performing foreground removal, we calculated the corresponding MAD values. In the test set for CMBFSCNN, the values of the MAD are K for recovery of CMB Q map and K for recovery of CMB U map. It can be observed that the MAD value obtained using the TCMB method is slightly smaller than that obtained using the CMBFSCNN method.
After noise debiasing, the recovered CMB power spectra is presented in Figure 12. It can be observed that the CMB BB power spectrum exhibits a slight deviation at large scales due to the influence of foreground residuals, while the impact on the recovery of the EE spectrum is relatively minimal. Based on the results from the recovered CMB sky map and power spectra, it can be concluded that the network model demonstrates a clear dependency on the foreground model. If there are significant changes in the foreground model, the effectiveness of foreground removal by the network model may be slightly diminished; however, the impact of these minor residual foregrounds on the results is limited. Deep learning methods are inherently data-driven algorithms. To mitigate the network’s dependence on foreground models, we can develop more diverse and comprehensive training datasets. For instance, the training set can incorporate a greater variety of samples from different foreground models. Additionally, we can generate new foreground maps by employing a weighted linear combination of sky maps produced by various foreground models. These strategies can enhance the diversity of the samples, thereby contributing to a reduction in the network’s reliance on specific foreground models. Consequently, investigating how to construct more appropriate datasets to support deep learning methodologies is of paramount importance. This work is reserved for future work.
4.5 Variation in the level of instrument noise
In actual observations, the level of instrumental noise could fluctuate. In this study, we investigate the impact of variations in instrumental noise levels on the results. After the network model has been trained on the standard training set, we re-simulate the testing dataset. Specifically, for the noise standard deviation maps of AliCPT at 95 GHz and 150 GHz, we multiply the value of each pixel by a random number drawn from a distribution with a mean of 1 and a variance of 0.1. In other words, each pixel in the noise standard deviation map is independently adjusted by 10%. We input the testing dataset, altered to reflect variations in noise levels, into the network model that has been trained on the standard training dataset, in order to evaluate the impact of noise level fluctuations on the model’s performance.
The reconstructed CMB maps are found to be largely consistent with the results presented in Figure 4. Here, we do not display specific results at map level; rather, we provide quantitative measures of the mean absolute deviation (MAD) to evaluate the effectiveness of foreground removal. The MAD value for the reconstructed CMB Q map is observed to be K, while the MAD for the U map is K. These MAD values are consistent with the results presented in Section 4.1. After the removal of noise bias, Figure 13 presents the reconstructed CMB polarization power spectra. It is evident that the reconstructed CMB polarization power spectrum is consistent with the results shown in Figure 6. These results suggest that minor variations in noise levels for each pixel have a negligible impact on the outcomes.
5 Conclusions
In this study, we propose a deep learning method based on the Transformer model, called TCMB, which is capable of effectively processing HEALPix spherical sky map. This method is applied to the foreground removal of CMB polarization in the ground based CMB polarization observations. The results demonstrate that our method can efficiently remove the CMB foreground, validating its effectiveness.
Based on the simulated data from the AliCPT-1 like experiment, we employed the TCMB method to perform foreground removal of CMB polarization. At the sky map level, the TCMB method effectively removes the foreground while retaining the CMB signal and instrument noise. We computed the cross-correlation between two half-mission maps to eliminate noise bias, and the results demonstrate that the TCMB method can effectively recover the CMB polarization signal, with the recovered EE and BB power spectra consistent with the true values.
We compared the TCMB method with the CMBFSCNN approach previously employed based on the CNN, and the results indicate that TCMB outperforms CMBFSCNN in recovering the CMB maps, while also overcoming the edge effects associated with the inability of CNN methods to directly process Healpix-format spherical sky maps. This demonstrates that the TCMB method has a distinct advantage over CNN approaches when dealing with Healpix-format spherical sky maps.
Additionally, we investigated the model dependency of TCMB, and the results indicate a clear dependence of TCMB on the foreground model. In scenarios where the foreground model undergoes significant changes, the efficiency of foreground removal may decrease; however, the impact of the residual foreground on the results is limited. Finally, we tested the effect of variations in instrument noise levels on the results, and the findings suggest that a 10% change in instrument noise levels has a negligible impact on the outcomes.
The TCMB method can effectively handle HEALPix spherical sky maps, making it applicable not only to the processing of CMB data but also to other experimental datasets, such as 21 cm signal data and galaxy survey data.
We thank Yi-Ming Wang, Yong-Ping Li, Hua Zhai, Wen-Zheng Chen for useful discussion. This study is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20241729, the China Postdoctoral Science Foundation under Grant Number 2024M763267, the National Nature Science Foundation of China under grant Nos. 12403005 12473004, and 12021003, the National Key R&D Program of China No. 2020YFC2201601 and 2020YFC2201603, and the Fundamental Research Funds for the Central Universities. Some of the results in this paper have been derived using the HEALPix (page: https://healpix.sourceforge.io/). We would like to express our gratitude for the following software packages: Healpy, CAMB, NaMaster, LensPyx, PySM, and PySM3.